A very small circular loop of radius r and carrying a current i1 is placed in the x-y plane with its centre on x-axis at the point C(a,0). A square loop of side length 2l carrying a current i2 is fixed in the y-z plane with the centre of the loop at the origin . Calculate the torque exerted by the square loop on the circular loop?

Grade:12

1 Answers

AskiitianExpert Shine
14 years ago

Hi

by symmetry we can conclude that the net magnetic feild due to the square loop at (a,o) is in the x direction
           So,
              B(net)=4*[(u0*i2/4*pi*x)*(2cosy)]*cos(90-y)
 
                where x=(a2+l2)1/2
                         y=tan-1(l/a)
                          
                 cos(90-y) is done as it gives x-component)

              

SO B=u0*i2*a*l/(pi*(x 3/2))

  now assuming B remains uniform throughout circular loop, torque=mxB
              where, m=magnetic moment of circular loop